OpenStudy (flor):

business and finance suppose that the cost c, in dollars of producing x chairs is given by, c= 2x^2 -40x+2400. how many chairs can be produce for P5,400

Parth (parthkohli):

Equate \( \color{Black}{\Rightarrow 2x^2 - 40x + 2400 = 5,400}\)

Parth (parthkohli):

Factor out the 2 for starting it. \( \color{Black}{\Rightarrow 2(x^2 - 20x + 1200) = 5,400}\) \( \color{Black}{\Rightarrow x^2 - 20x + 1200 = 2,700}\)

Parth (parthkohli):

Can you solve the above equation?

OpenStudy (flor):

can you give the representation??????????

Parth (parthkohli):

Representation?